### v4

@@ Appendix A Beyond Volatility Factors
Appendix B Simulation Study
We present a simulation exercise to illustrate how the CIQ premiums would look like if the driving force behind them were simply common volatility. We simulate stock returns from the following model
@@ Appendix B Simulation Study
Time-series variation of the returns drive two common factors – market factor, , and variance factor . The expected return of a stock is then equal to
@@
where and are the previous month’s log-variance and average log-variance over the last 12-month period, respectively. The common variance process is approximated by the cross-sectional average of the squared residuals from the time series regression of stock returns on the market factor. We fit the model from equation 22 on this time series. When simulating this time series, we initialise the process by randomly selecting 12 consequent observations of the common variance process estimated from the data and using those observations for iterating forward.
We calibrate the simulation setting to match the CRSP data sample we employ in the empirical investigation. We estimate stock-level market beta, , using time-series regression of stock return on the market return. Exposure to the common variance, , is estimated by regressing the stock return on the estimate of the common variance process. Price of risk associated with the variance exposure, is chosen to be equal to .282828This value corresponds to approximately 6% annual high minus low premium obtained from ten portfolios portfolios sorted on the exposure to the common variance. The choice of this value is not essential for the results that we present here. We estimate stock-level parameters of the idiosyncratic error variance–, –as the sample mean and standard deviation of . To approximate the , we use squared residuals from the time-series regression of the stock return on the market return. Then, to simulate these parameters, we approximate their distribution by normal distribution, with the mean equal to the estimates’ cross-sectional average and the variance equal to the cross-sectional variance of the estimates.
@@
Similarly, as in the empirical investigation, we sort stocks into decile portfolios based on their estimated exposure to the factors to infer the associated risk premiums. We proxy the premiums by computing high-minus-low returns of the portfolios. Table 10 reports the average premiums for the three CIQ factors that we investigate in the empirical analysis. We observe that the premium is positive for the lower-tail quantile factor, negative for the upper-tail factor and close to zero for the central factor. The magnitude of the premiums are comparable across lower and upper tail factors and in absolute value approximately equal to 9.4%. The premium associated with the exposure to the PCA-SQ factor is -6.09%. We also compute associated -statistics as a ratio between average premium and its standard deviation across all the simulations. Premiums for the lower and upper tail factors are significant, unlike the value for the central factor, with -statistics of around 2.6 in absolute value. The -value associated with the PCA-SQ factor is -2.33, so the premium estimated using this approach is also significant. Next, we present the proportion of rejections of non-significance of the CIQ betas at a 5% significance level from multivariate cross-sectional regressions of average returns on those betas and market betas. We can see that the proportions are virtually identical for both lower- and upper-tail betas of around 96%. The ratio for the PCA-SQ betas is 90%.
The table contains risk average premiums computed from high-minus-low returns of decile portfolios sorted on exposure to the CIQ() risks. We simulate the returns using common variance factor model proposed by Ding et al. (2025). We simulate panel of 2,500 stocks with 120 monthly observations. We perform the simulation 1,000 times. -statistics are obtained by dividing the average premium by its standard deviation. We also report proportion of rejections of non-significance of CIQ betas from multivariate cross-sectional regressions of average returns on those betas and market betas.
@@
As we can see from the results, if there was a simple common volatility element present in the returns, which is compensated in the cross-section, the CIQ risk premium would be symmetrical for lower and upper tail quantile factors. Moreover, the exposure to the PCA-SQ factor would be priced in this case, too. Overall, the evidence from the simulation exercise suggests that the CIQ risk premiums we observe in the data are not attributable to the common volatility compensation.
Appendix C Cross-Sectional Quantiles
The table provides a summary of cross-sectional dispersion of excess stock returns. Each month, we compute cross-sectional quantile of raw or idiosyncratic excess stock returns with respect to market factor (CAPM), three factors of Fama and French (1993), five factors of Fama and French (2015) (FF5) or six factors of Fama and French (2018) (FF6). We report results for 20% (lower-tail) and 80% (upper-tail) quantiles. In Panel A, we report variance ratios between time-series of these quantile series. Rows correspond to the quantile series in the nominator, while columns corresponds to the quantile series in the denominator. In Panel B, we report regression results from regressing FF3 idiosyncratic quantiles on intercept, idiosyncratic volatility and CIQ (either lower- or upper-tail) factor. The data come from the CRSP database and cover the period from January 1963 to December 2018.
@@ Appendix C Cross-Sectional Quantiles
| Intercept | -0.09 | -0.05 | 0.00 | 0.08 | 0.03 | 0.01 |
| (-101.89) | (-22.68) | (-1.06) | (83.60) | (14.32) | (3.17) | |
| Volatility | -0.24 | 0.29 | ||||
| (-17.39) | (21.60) | |||||
| CIQ factor | 0.08 | 0.07 | ||||
@@
| adj. | 0.00 | 0.32 | 0.49 | 0.00 | 0.42 | 0.46 |
The figure shows various measures of cross-sectional stock returns dispersion through time. For each month, we plot 20% and 80% cross-sectional quantiles of either excess or idiosyncratic excess returns. The idiosyncratic returns are computed with respect to the three-factor model of Fama and French (1993). The data come from the CRSP database and cover the period from January 1963 to December 2018. The shaded areas represent NBER recessions.
Appendix D Risk Measures Definitions
@@ D.1 Market Beta
D.2 Idiosyncratic Volatility
D.3 Total and Idiosyncratic Skewness
@@ D.6 PCA-SQ Betas
D.7 CIV Beta
CIV beta is estimated by regressing monthly excess stock returns on a constant and increments of the CIV factor of Herskovic et al. (2016). We use 60-month rolling window and require at least 48 observations.
D.8 VIX Beta
@@ D.9 Downside Beta
D.10 Tail Risk Beta
@@
D.11 Hybrid Tail Covariance Risk
@@
D.12 Multivariate Crash Risk
@@
D.13 Predicted Systematic Co-skewness
@@
where is the co-skewness from Equation 36 computed using daily returns from month to month , are risk measures (volatility, market beta, etc.) estimated using daily data from month to month , and are characteristics (size, book-to-price, etc.) observed at the end of month . The function transforms the original variable into its normalised rank in the cross-section of variable , which posses observations.
@@
Appendix E Summary of the CIQ Factors

### v5

@@ Appendix B Beyond Volatility Factors
Appendix C A Simple Intermediary-Fragility Model
This appendix provides a simple model that formalizes the mechanism in Section 3.2. The goal is not to offer a fully calibrated structural model, but to show that a continuum of intermediary balance-sheet constraints can aggregate into a one-dimensional fragility state that (i) drives common shifts in firms’ idiosyncratic left-tail quantiles and (ii) appears in the stochastic discount factor through innovations .
C.1 Setup
There is a continuum of competitive intermediaries with equity . Intermediary chooses a portfolio over a market factor and a continuum of firm claims with returns , subject to an expected shortfall (ES) constraint on portfolio losses. Let denote the Lagrange multiplier on this constraint.
We define systemic fragility as the equity-weighted average multiplier
where are equity weights. Intuitively, is the aggregate shadow cost of risk-bearing in the intermediary sector.
Firm ’s idiosyncratic return has a conditional quantile representation
with for , for , and satisfies the appropriate quantile restriction. This captures the idea that tighter constraints shift firms’ left-tail quantiles downward through funding, covenant, and liquidity channels, but leave the center and right tail largely unaffected.
C.2 Quantile factor representation
Proposition 2 (Quantile factor representation).
Under standard large- conditions for approximate factor models, the first quantile principal component of idiosyncratic returns at consistently estimates an affine transformation of :
with and an error term .
Sketch of proof. Equation ((27)) defines a one-factor structure in the conditional -quantiles of idiosyncratic returns. Under the assumptions of Chen et al. (2021) on cross-sectional pervasiveness and bounded idiosyncratic components, the quantile factor analysis estimator recovers the space spanned by up to an affine transformation. Normalization conditions pin down the scale and sign up to a constant.
C.3 Pricing kernel
Each intermediary’s optimality condition with an ES constraint yields an SDF containing a term proportional to its marginal shortfall loss , in addition to a standard component . In equilibrium, all intermediaries share the same pricing kernel, which can be written as
where is the aggregate marginal shortfall and is proportional to the cross-sectional average of , hence to by (26).
Assuming that portfolio shortfalls materialize precisely when the fragility state worsens, we can approximate as a linear function of :
with . Substituting into (28) yields
with in constrained states.
Proposition 3 (CIQ beta as fragility exposure).
With , the innovation is proportional to up to noise. The CIQ beta
is therefore proportional to the covariance of with . Expected excess returns satisfy
with when constraints are tight.
Sketch of proof. Combine with (29) and apply the standard linear beta pricing argument. Under mild assumptions on the noise term , the projection of on coincides with its projection on up to a constant.
Proposition 4 (Left-tail asymmetry).
Because ES constraints depend on losses, the marginal shortfall is insensitive to upside co-movement and depends only on left-tail realizations. Under the quantile representation (27), the resulting price of quantile-factor risk is monotone in for and negligible for . In particular, only the lower-tail CIQ factor is priced, while median and upper-tail CIQ factors do not command premia.
Sketch of proof. The ES constraint binds only in states where portfolio returns fall below a left-tail threshold. By construction, is orthogonal to purely central or right-tail shifts. Under (27), affects only for . Thus, only quantile factors estimated from the left tail proxy for and receive a non-zero price of risk in (29).
Appendix D Simulation Study
We present a simulation exercise to illustrate how the CIQ premia would look like if the driving force behind them were simply common volatility. We simulate stock returns from the following model
@@ Appendix D Simulation Study
Time-series variation of the returns drive two common factors – market factor, , and variance factor with unconditional mean . The expected return of a stock is then equal to
@@
where and are the previous month’s log-variance and average log-variance over the last 12-month period, respectively. The common variance process is approximated by the cross-sectional average of the squared residuals from the time series regression of stock returns on the market factor. We fit the model from Equation (33) on this time series. When simulating this time series, we initialise the process by randomly selecting 12 consequent observations of the common variance process estimated from the data and using those observations for iterating forward.
We calibrate the simulation setting to match the CRSP data sample we employ in the empirical investigation. We estimate stock-level market beta, , using time-series regression of stock return on the market return. Exposure to the common variance, , is estimated by regressing the stock return on the estimate of the common variance process. Price of risk associated with the variance exposure, is chosen to be equal to .262626This value corresponds to approximately 6% annual high minus low premium obtained from ten portfolios portfolios sorted on the exposure to the common variance. The choice of this value is not essential for the results that we present here. We estimate stock-level parameters of the idiosyncratic error variance–, –as the sample mean and standard deviation of . To approximate the , we use squared residuals from the time-series regression of the stock return on the market return. Then, to simulate these parameters, we approximate their distribution by normal distribution, with the mean equal to the estimates’ cross-sectional average and the variance equal to the cross-sectional variance of the estimates.
@@
Similarly, as in the empirical investigation, we sort stocks into decile portfolios based on their estimated exposure to the factors to infer the associated risk premia. We proxy the premia by computing high-minus-low returns of the portfolios. Table 13 reports the average premia for the three CIQ factors that we investigate in the empirical analysis. We observe that the premium is positive for the lower-tail quantile factor, negative for the upper-tail factor and close to zero for the central factor. The magnitude of the premia are comparable across lower and upper tail factors and in absolute value approximately equal to 9.4%. The premium associated with the exposure to the PCA-SQ factor is -6.09%. We also compute associated -statistics as a ratio between average premium and its standard deviation across all the simulations. premia for the lower and upper tail factors are significant, unlike the value for the central factor, with -statistics of around 2.6 in absolute value. The -value associated with the PCA-SQ factor is -2.33, so the premium estimated using this approach is also significant. Next, we present the proportion of rejections of non-significance of the CIQ betas at a 5% significance level from multivariate cross-sectional regressions of average returns on those betas and market betas. We can see that the proportions are virtually identical for both lower- and upper-tail betas of around 96%. The ratio for the PCA-SQ betas is 90%.
The table contains risk average premia computed from high-minus-low returns of decile portfolios sorted on exposure to the CIQ() risks. We simulate the returns using common variance factor model proposed by Ding et al. (2025). We simulate panel of 2,500 stocks with 120 monthly observations. We perform the simulation 1,000 times. -statistics are obtained by dividing the average premium by its standard deviation. We also report proportion of rejections of non-significance of CIQ betas from multivariate cross-sectional regressions of average returns on those betas and market betas.
@@
As we can see from the results, if there was a simple common volatility element present in the returns, which is compensated in the cross-section, the CIQ risk premium would be symmetrical for lower and upper tail quantile factors. Moreover, the exposure to the PCA-SQ factor would be priced in this case, too. Overall, the evidence from the simulation exercise suggests that the CIQ risk premia we observe in the data are not attributable to the common volatility compensation.
Appendix E Cross-Sectional Quantiles
The table provides a summary of cross-sectional dispersion of excess stock returns. Each month, we compute cross-sectional quantile of raw or idiosyncratic excess stock returns with respect to market factor (CAPM), three factors of Fama and French (1993), five factors of Fama and French (2015) (FF5) or six factors of Fama and French (2018) (FF6). We report results for 20% (lower-tail) and 80% (upper-tail) quantiles. In Panel A, we report variance ratios between time-series of these quantile series. Rows correspond to the quantile series in the denominator, while columns corresponds to the quantile series in the nominator. In Panel B, we report regression results from regressing FF3 idiosyncratic quantiles on intercept, idiosyncratic volatility and CIQ (either lower- or upper-tail) factor. The data come from the CRSP database and cover the period from January 1968 to December 2024.
@@ Appendix E Cross-Sectional Quantiles
| Intercept | -0.09 | -0.05 | -0.02 | 0.08 | 0.03 | 0.02 |
| (-106.88) | (-23.72) | (-4.35) | (88.70) | (15.96) | (5.37) | |
| Volatility | -0.23 | 0.26 | ||||
| (-17.83) | (20.48) | |||||
| CIQ factor | 0.07 | 0.06 | ||||
@@
| adj. | 0.00 | 0.32 | 0.40 | 0.00 | 0.38 | 0.39 |
The figure shows various measures of cross-sectional stock returns dispersion through time. For each month, we plot 20% and 80% cross-sectional quantiles of either excess or idiosyncratic excess returns. The idiosyncratic returns are computed with respect to the three-factor model of Fama and French (1993). The data come from the CRSP database and cover the period from January 1968 to December 2024. The shaded areas represent NBER recessions.
Appendix F Risk Measures Definitions
@@ F.1 Market Beta
F.2 Idiosyncratic Volatility
F.3 Total and Idiosyncratic Skewness
@@ F.6 PCA-SQ Betas
F.7 CIV Beta
Following Herskovic et al. (2016), CIV beta is estimated by regressing monthly excess stock returns on a constant, increments of the CIV factor and increments of the monthly market variance
We use 60-month rolling window and require at least 48 observations.
F.8 VIX Beta
@@ F.9 Downside Beta
F.10 Tail Risk Beta
@@
F.11 Hybrid Tail Covariance Risk
@@
F.12 Multivariate Crash Risk
@@
F.13 Predicted Systematic Co-skewness
@@
where is the co-skewness from Equation (48) computed using daily returns from month to month , are risk measures (volatility, market beta, etc.) estimated using daily data from month to month , and are characteristics (size, book-to-price, etc.) observed at the end of month . The function transforms the original variable into its normalised rank in the cross-section of variable , which posses observations.
@@
Appendix G Summary of the CIQ Factors